### v1

@@ IV-B Optimal Policy
Before presenting the optimal policy, we define the index function. For every $\theta\in\mathbb{N}$, the index function of modality $m$, denoted by $\gamma_{m}$, is defined as
$\gamma_{m}(\theta)\coloneq\inf_{k\in\mathbb{Z}_{+}}\frac{C_{m}(\theta+k)-C_{m}(\theta)}{kT_{m}},$ (4)
where the numerator is the additional total inference error and the denominator is the additional transition duration, both when $\tau_{m}$ is $\theta+k$ instead of $\theta$ (see Remark 1 in Section IV-D for a detailed explanation of the index function).
@@
Suppose $L$ is uniformly bounded (Assumption 1). For every $\beta\in\mathbb{R}$, define
$\tau_{m,\mathrm{opt}}(\beta)=\inf\{\theta\in\mathbb{N}:\gamma_{m}(\theta)\geq\beta\},\quad\forall m\in\{1,2\},$ (5)
where the index function $\gamma_{m}$ of each modality $m$ is defined in Eq. (4). An optimal policy for Problem OPT, represented by $\tau_{m,\mathrm{opt}}$ for each $m\in\{1,2\}$, is then given by
$\tau_{m,\mathrm{opt}}=\tau_{m,\mathrm{opt}}(\bar{L}_{\mathrm{opt}})=\inf\{\theta\in\mathbb{N}:\gamma_{m}(\theta)\geq\bar{L}_{\mathrm{opt}}\},$ (6)
@@
$g_{2}(\beta)$ $\coloneq\left(\tau_{1,\mathrm{opt}}(\beta)+1\right)T_{1}+\left(\tau_{2,\mathrm{opt}}(\beta)+1\right)T_{2}.$
The threshold $\bar{L}_{\mathrm{opt}}$ is a root of
$g_{1}(\beta)-\beta g_{2}(\beta)=0.$ (7)
We will prove the theorem in Section IV-C. Eq. (5) suggests that the optimal policy exhibits an index-based threshold structure. That is, the policy selects modality $m$ for $\theta$ consecutive transmissions, stopping at the first time when the index function $\gamma_{m}(\theta)$ exceeds a threshold $\beta$. Then the policy switches to select the other modality. The index function for each modality can be pre-computed independently, as it depends only on the known parameters (i.e., the inference error function and the transmission times).
Moreover, Eq. (6) shows that, under the optimal policy, both modalities share the same threshold $\bar{L}_{\mathrm{opt}}$, which is exactly the optimal objective value of Problem OPT. The threshold $\bar{L}_{\mathrm{opt}}$ can be determined by solving Eq. (7). As will be shown later, Eq. (7) is obtained by substituting $\beta$ and $\tau_{m,\mathrm{opt}}(\beta)$ into the Bellman optimality equation (3). The following proposition shows that Eq. (7) can be efficiently solved.
@@
Define
$g(\beta)\coloneq g_{1}(\beta)-\beta g_{2}(\beta).$ (8)
@@
The function $g$ is the left-hand side of Eq. (7). It can be shown that $g$ is essentially the infimum of multiple linear functions of $\beta$ with negative coefficients, which leads to the properties stated above. The complete proof is similar to that of [1, Lemma 9] and is therefore omitted. ∎
@@ IV-C Proof of Theorem 1
We prove Theorem 1 by showing that our policy satisfies the Bellman optimality equation (3) and the optimal solution $\tau_{m,\mathrm{opt}}$ attains the infimum in Eq. (3) for every $m$. According to average-cost MDP theory [19, Proposition 4.2.1], if there exists a finite scalar $\bar{L}_{\mathrm{opt}}$ and a uniformly bounded function $h$ that solves the Bellman optimality equation, then an optimal policy exists. Furthermore, a policy is optimal if it attains the infimum in the Bellman optimality equation for all states.
We cannot directly solve each equation independently, because the optimal objective value $\bar{L}_{\mathrm{opt}}$ appears in the Bellman optimality equation (3) for each state $\mathbf{\Delta}_{m,\mathrm{re}}$. To this end, we fix $\bar{L}_{\mathrm{opt}}=\beta$ for arbitrary $\beta\in\mathbb{R}$, and focus on solving Problem OPT-$\beta$, defined as
$\textbf{OPT-$\beta$}\quad\inf_{\tau\in\mathbb{N}}\left[C_{m}(\tau)-\tau T_{m}\beta\right],\quad\forall m\in\{1,2\}.$
From the Bellman optimality equation (3) to Problem OPT-$\beta$, we discard two terms: $T_{m^{\prime}}\bar{L}_{\mathrm{opt}}$ and $h(\mathbf{\Delta}_{m^{\prime},\mathrm{re}})$, as they are independent of the decision $\tau_{m}$. Note that Problem OPT-$\beta$ can be solved separately for each modality. The next proposition shows that the optimal solution for Problem OPT-$\beta$ can be expressed in terms of $\beta$.
@@
Suppose $L$ is uniformly bounded (Assumption 1). For every given $\beta$ and $m$, the optimal solution to Problem OPT-$\beta$ is given by $\tau_{m,\mathrm{opt}}(\beta)$ in Eq. (5), i.e.,
$\tau_{m,\mathrm{opt}}(\beta)=\inf\left\{\theta\in\mathbb{N}:\gamma_{m}(\theta)\geq\beta\right\},$
where the index function $\gamma_{m}$ is defined in Eq. (4).
@@
We use induction to show that $\tau_{m,\mathrm{opt}}(\beta)=\theta$ attains the infimum of Problem OPT-$\beta$ if it satisfies Eq. (5). Firstly, we show that $\tau_{m,\mathrm{opt}}(\beta)=0$ if $\gamma_{m}(0)\geq\beta$. Then, assume that the $\tau_{m,\mathrm{opt}}(\beta)$ satisfies Eq. (5) when $\tau_{m,\mathrm{opt}}(\beta)=0,1,\dots,\theta-1$. We can show that $\tau_{m,\mathrm{opt}}(\beta)=\theta$ if (i) $\gamma_{m}(i)<\beta$ for $i=0,1,\dots,\theta-1$ and (ii) $\gamma_{m}(\theta)\geq\beta$. Combining conditions (i) and (ii), we derive Eq. (5). See the Appendix for the detailed proof. ∎
Proposition 2 shows that the optimal solution $\tau_{m,\mathrm{opt}}(\beta)$ for Problem OPT-$\beta$ attains the infimum in the Bellman optimality equation (3) when $\bar{L}_{\mathrm{opt}}=\beta$. Hence, $\tau_{m,\mathrm{opt}}(\bar{L}_{\mathrm{opt}})$ given in Eq. (6) is an optimal policy for Problem OPT.
Next, we show that $\bar{L}_{\mathrm{opt}}$ is a root of Eq. (7) by constructing a solution to the Bellman optimality equation (3). We substitute $\bar{L}_{\mathrm{opt}}=\beta$ and $\tau_{m,\mathrm{opt}}(\beta)$ into the Bellman optimality equation (3) and obtain
@@
By summing the two equations, canceling $h(\mathbf{\Delta}_{m,\mathrm{re}})$, and rearranging the terms, we obtain a function of $\beta$:
$C_{m}(\tau_{m^{\prime},\mathrm{opt}}(\beta))+C_{m^{\prime}}(\tau_{m^{\prime},\mathrm{opt}}(\beta))$
@@
which is exactly Eq. (7) in Theorem 1. Finally, we show that $\bar{L}_{\mathrm{opt}}$ and $h(\mathbf{\Delta}_{m,\mathrm{re}})$ exist and are bounded, which concludes the proof of Theorem 1. With the two properties stated in Proposition 2 and by applying the Intermediate Value Theorem [21], we conclude that Eq. (7) has a unique and finite root. Thus, $\bar{L}_{\mathrm{opt}}$ exists and is finite. By substituting $\tau_{m,\mathrm{opt}}(\bar{L}_{\mathrm{opt}})$ and $h(\mathbf{\Delta}_{m^{\prime},\mathrm{re}})=0$ into the Bellman optimality equation, we obtain $h(\mathbf{\Delta}_{m,\mathrm{re}})$. Furthermore, $h(\mathbf{\Delta}_{m,\mathrm{re}})$ is bounded because $\bar{L}_{\mathrm{opt}}$, $\tau_{m,\mathrm{opt}}$, $T_{m}$, and $L$ are bounded. ∎
@@ IV-D Discussions
The index function $\gamma_{m}(\theta)$ reflects the minimum future cost if the scheduler continues to select modality $m$ after having selected it for $\theta$ consecutive transmissions. To see this, consider a symmetric case when $T_{1}=T_{2}=T_{c}$ for some constant $T_{c}\in\mathbb{Z}_{+}$. Then, the index function for modality 1 in Eq. (5) becomes
$\gamma_{1}(\theta)=\inf_{k\in\mathbb{Z}_{+}}\frac{\sum_{j=1}^{k}\sum_{i=0}^{T_{c}-1}L(T_{c}+i,(\theta+2+j)T_{c}+i)}{kT_{c}},$
@@
$C_{b}(\theta,k)\coloneq$ $\sum_{i=0}^{T_{2}-1}L(T_{1}+i,(\theta+k+1)T_{1}+T_{2}+i)$
$+\sum_{i=T_{2}}^{T_{1}-1}L(T_{1}+i,(\theta+2)T_{2}+i).$
The index function for modality 1 then becomes
$\gamma_{1}(\theta)=\inf_{k\in\mathbb{Z}_{+}}\frac{C_{a}(\theta,k)+C_{b}(\theta,k)}{kT_{1}},$
where $C_{a}(\theta,k)$ quantifies the total future cost and $C_{b}(\theta,k)$ adjusts the cost due to different transmission times.
@@
Our results generalize the two-source scheduling in remote estimation [22]. Specifically, when the inference error function $L$ is non-decreasing and the two modalities have unit transmission times (i.e., $T_{1}=T_{2}=1$), the index function of modality 1 reduces to
$\gamma_{1}(\theta)=\inf_{k\in\mathbb{Z}_{+}}\frac{\sum_{j=\theta+2}^{\theta+k+1}L(1,j+1)}{k}=L(1,\theta+3),$
where the last equality holds as the infimum is achieved when $k=1$ as $L$ is non-decreasing. Then, the optimal policy for modality 1 is $\tau_{1,\mathrm{opt}}=\inf\{\theta:L(1,\theta+3)\geq\bar{L}_{\mathrm{opt}}\}$, which can be shown to be equivalent to the result in [22, Proposition 3.5]. Note that in this special case, the index function is non-decreasing, whereas in our setting, it can be non-monotonic.
@@
A similar index-based threshold policy has been studied in the single-source scheduling for remote inference [1, 13]. Specifically, under stationary random transmission time, the optimal time to send a new feature is when an AoI-based index function exceeds a threshold. In contrast, we apply the index function to the optimal modality scheduling problem, which differs from the single-source setting.
@@ V Numerical Case Study: Robot State Prediction
Figure 4: Inference error vs. AoI.
@@
Figure 5: Performance comparison under varying transmission times.
V-A System Description
Consider a robot performing tasks such as lifting objects or walking. It gathers both environmental and self-state information using multimodal sensors, such as LiDAR, cameras, and onboard sensors. On the receiver side, a predictor aims to continuously track the robot’s state, such as its pose and velocity. A transmitter sends the features collected from the robot to the receiver, where an ML model is used to predict the robot’s state based on the received data. Due to the high dimensionality and varying sizes of the data, transmission often spans multiple time slots and may differ across modalities. Next, we describe our experimental setup.
V-B Experimental Setup
We consider the OpenAI Bipedal Walker as our robot task, where a four-joint robot must run over stumps, pitfalls, and other obstacles (see [23] for details). The reinforcement learning algorithm used to control the robot is TD3-FORK [24], which achieves state-of-the-art performance on this task. After training the control model, we generate a time-series dataset in the OpenAI Gymnasium simulation environment, consisting of LiDAR rangefinder measurements, robot state information, and joint control signals.
For the state prediction task, we aim to predict the robot’s joint velocities using sequential LiDAR measurements and joint control signals as two distinct modalities. We adopt the Long Short-Term Memory (LSTM) neural network as the predictor model, due to its widespread use and effectiveness in time-series forecasting. The network architecture includes an input layer, a hidden layer with 20 LSTM cells, and a fully connected output layer. We use 80% of the dataset for training. To incorporate AoI into model training, we augment the dataset as follows: given any AoI vector $(\delta_{1},\delta_{2})$, we construct the feature-label pairs $(X_{1,t-\delta_{1}},X_{2,t-\delta_{2}};Y_{t})$ for all data points, where the input features are aligned with their corresponding AoI values. The LSTM network takes the AoI vector as two additional input features and is trained on the augmented training dataset.

### v2

@@ IV-B Optimal Policy
Before presenting the optimal policy, we define the index function. For every $\theta=0,1,\dots,\tau_{\max}-1$, the index function of modality $m$, denoted by $\gamma_{m}(\theta)$, is defined as
$\gamma_{m}(\theta)\coloneq\min_{k\in\{1,2,\dots,\tau_{\max}-\theta\}}\frac{C_{m}(\theta+k)-C_{m}(\theta)}{kT_{m}},$ (4)
where the numerator is the additional total inference error and the denominator is the additional transition duration, both when the decision is $\theta+k$ instead of $\theta$ (see Remark 1 in Section IV-D for a detailed explanation of the index function).
@@
Suppose $L$ is uniformly bounded (Assumption 1), $\inf\emptyset=\infty$, and $\theta\in\{0,1,\dots,\tau_{\max}-1\}$. For every $\beta\in\mathbb{R}$ and $m\in\{1,2\}$, define
$\tau_{m,\mathrm{opt}}(\beta)\coloneq\min\big{\{}\min\{\theta:\gamma_{m}(\theta)\geq\beta\},\tau_{\max}\big{\}}.$ (5)
The index function $\gamma_{m}$ of each modality $m$ is defined in Eq. (4). The optimal stationary policy for Problem OPT, represented by $(\tau^{*}_{1},\tau^{*}_{2})$, is then given by
$\tau^{*}_{m}=\tau_{m,\mathrm{opt}}(\bar{L}_{\mathrm{opt}}),\quad\forall m\in\{1,2\},$ (6)
@@
$g_{2}(\beta)$ $\coloneq\left(\tau_{1,\mathrm{opt}}(\beta)+1\right)T_{1}+\left(\tau_{2,\mathrm{opt}}(\beta)+1\right)T_{2},$
and let $g(\beta)\coloneq g_{1}(\beta)-\beta g_{2}(\beta)$. The threshold $\bar{L}_{\mathrm{opt}}$ is the unique root of the following equation:
$g(\beta)=0.$ (7)
We will prove Theorem 1 in Section IV-C. Eq. (5) in Theorem 1 suggests that the optimal stationary policy exhibits an index-based threshold structure. That is, the policy selects modality $m$ for $\theta$ consecutive transmissions, stopping at the first time when the index function $\gamma_{m}(\theta)$ exceeds a threshold $\beta$; then, the policy switches to select the other modality. The index function of each modality can be pre-computed independently, as it depends only on the known parameters (i.e., the inference error function and the transmission times).
Moreover, Eq. (6) shows that, under the optimal policy, both modalities share the same threshold $\bar{L}_{\mathrm{opt}}$, which is exactly the optimal objective value of Problem OPT. The threshold $\bar{L}_{\mathrm{opt}}$ can be determined by solving Eq. (7). The following proposition shows that Eq. (7) can be efficiently solved.
@@
It turns out that $g$ is the minimum of multiple linear functions of $\beta$ with negative coefficients, which leads to the stated properties. The complete proof is similar to that of [1, Lemma 9] and is therefore omitted. ∎
@@ IV-C Proof of Theorem 1
We prove Theorem 1 by showing that our policy satisfies the Bellman optimality equation (3) and the optimal decision $(\tau^{*}_{1},\tau^{*}_{2})$ attains the minimum in Eq. (3). According to average-cost SMDP theory [19, Theorem 11.4.4] [21], for an SMDP with countable state and decision space, if a finite scalar $\bar{L}_{\mathrm{opt}}$ and a uniformly bounded function $h$ satisfy the Bellman optimality equation, then an optimal stationary policy exists. Furthermore, a policy is optimal if it attains the minimum in the Bellman optimality equation for all states.
We cannot directly solve each Bellman optimality equation (3) independently, because the optimal objective value $\bar{L}_{\mathrm{opt}}$ appears in the Bellman optimality equation (3) for each state $\mathbf{\Delta}_{m,\mathrm{re}}$. To solve this, we fix $\bar{L}_{\mathrm{opt}}=\beta$ for arbitrary $\beta\in\mathbb{R}$ and focus on solving Problem OPT-$\beta$, defined as
$\textbf{OPT-$\beta$}\quad\min_{\tau\in\{0,1,\dots,\tau_{\max}\}}\left[C_{m}(\tau)-\tau T_{m}\beta\right],\quad\forall m\in\{1,2\}.$
From the Bellman optimality equation (3) to Problem OPT-$\beta$, we discard two terms: $T_{m^{\prime}}\bar{L}_{\mathrm{opt}}$ and $h(\mathbf{\Delta}_{m^{\prime},\mathrm{re}})$, as they are independent of the decision $\tau$. Note that Problem OPT-$\beta$ can be solved separately for each modality. The next proposition shows that the optimal solution for Problem OPT-$\beta$ can be expressed in terms of $\beta$.
@@
Suppose $L$ is uniformly bounded (Assumption 1), $\inf\emptyset=\infty$, and $\theta\in\{0,1,\dots,\tau_{\max}-1\}$. For every $\beta\in\mathbb{R}$ and $m\in\{1,2\}$, the optimal solution to Problem OPT-$\beta$ is given by $\tau_{m,\mathrm{opt}}(\beta)$ in Eq. (5), i.e.,
$\tau_{m,\mathrm{opt}}(\beta)\coloneq\min\big{\{}\min\{\theta:\gamma_{m}(\theta)\geq\beta\},\tau_{\max}\big{\}}.$
The index function $\gamma_{m}$ of each modality $m$ is defined in Eq. (4).
@@
Suppose $\tau_{m,\mathrm{opt}}(\beta)$ attains the minimum of Problem OPT-$\beta$. We use induction to show that for every integer $0\leq i\leq\tau_{\max}$, if $i$ satisfies Eq. (5), i.e., $\min\big{\{}\min\{\theta:\gamma_{m}(\theta)\geq\beta\},\tau_{\max}\big{\}}=i$, then $\tau_{m,\mathrm{opt}}(\beta)=i$. Firstly, we show that $\tau_{m,\mathrm{opt}}(\beta)=0$ if $\gamma_{m}(0)\geq\beta$. Assume that for any integer $1\leq j\leq\tau_{\max}$, the result holds for $i=0,1,\dots,j-1$, we aim to show the result holds for $i=j$. If $j<\tau_{\max}$, we show that $\tau_{m,\mathrm{opt}}(\beta)=j$ if (i) $\gamma_{m}(j)<\beta$ for $i=0,1,\dots,j-1$ and (ii) $\gamma_{m}(j)\geq\beta$; that is, $\tau_{m,\mathrm{opt}}(\beta)=j=\min\{\theta:\gamma_{m}(\theta)\geq\beta\}$. If $j=\tau_{\max}$, we show that $\tau_{m,\mathrm{opt}}(\beta)=\tau_{\max}=j$. Combining the two cases yields $\tau_{m,\mathrm{opt}}(\beta)=j$. See the Appendix for the detailed proof. ∎
Proposition 2 shows that the optimal solution $\tau_{m,\mathrm{opt}}(\beta)$ attains the minimum of the Bellman optimality equation (3) if $\bar{L}_{\mathrm{opt}}=\beta$. Hence, $\tau_{m,\mathrm{opt}}(\bar{L}_{\mathrm{opt}})$ given in Eq. (6) is an optimal policy for Problem OPT.
Next, we show that $\bar{L}_{\mathrm{opt}}$ is the unique root of Eq. (7) by constructing a solution to the Bellman optimality equation (3). Given any $\beta\in\mathbb{R}$, we substitute $\bar{L}_{\mathrm{opt}}=\beta$ and $\tau_{m,\mathrm{opt}}(\beta)$ into the Bellman optimality equation (3) and obtain
@@
By summing the two equations, canceling $h(\mathbf{\Delta}_{m,\mathrm{re}})$, and rearranging terms, we obtain an equation of $\beta$:
$C_{m}(\tau_{m,\mathrm{opt}}(\beta))+C_{m^{\prime}}(\tau_{m^{\prime},\mathrm{opt}}(\beta))$
@@
which is exactly Eq. (7) in Theorem 1. Finally, we conclude the proof by showing that $\bar{L}_{\mathrm{opt}}$ and $h(\mathbf{\Delta}_{m,\mathrm{re}})$ exist and are bounded. Using Proposition 1 and the Intermediate Value Theorem [22], it follows that Eq. (7) has a unique, finite root, i.e., a finite $\bar{L}_{\mathrm{opt}}$ exists. Hence, the optimal decision $\tau_{m,\mathrm{opt}}(\bar{L}_{\mathrm{opt}})$ in Eq. (5) exists, and trivially, its value is bounded by $\tau_{\max}$. Substituting $h(\mathbf{\Delta}_{m^{\prime},\mathrm{re}})=0$ and $\tau_{m}^{*}=\tau_{m,\mathrm{opt}}(\bar{L}_{\mathrm{opt}})$ into the Bellman optimality equation (3) yields
$h(\mathbf{\Delta}_{m,\mathrm{re}})=C_{m}(\tau_{m}^{*})-(\tau_{m}^{*}T_{m}+T_{m^{\prime}})\bar{L}_{\mathrm{opt}}.$
Finally, $h(\mathbf{\Delta}_{m,\mathrm{re}})$ is bounded because $\bar{L}_{\mathrm{opt}}$, $\tau_{m}^{*}$, $T_{m}$, and $L$ are all bounded. ∎
@@ IV-D Discussions
The index function $\gamma_{m}(\theta)$ reflects the minimum future cost if the scheduler continues to select modality $m$ after having selected it for $\theta$ consecutive transmissions. To see this, consider a symmetric case when $T_{1}=T_{2}=T_{c}$ for some constant $T_{c}\in\mathbb{Z}_{+}$. Then, the index function of modality 1 in Eq. (5) becomes
$\gamma_{1}(\theta)=\min_{k}\frac{\sum_{j=1}^{k}\sum_{i=0}^{T_{c}-1}L(T_{c}+i,(\theta+2+j)T_{c}+i)}{kT_{c}},$
@@
$C_{b}(\theta,k)\coloneq$ $\sum_{i=0}^{T_{2}-1}L(T_{1}+i,(\theta+1+k)T_{1}+T_{2}+i)$
$+\sum_{i=T_{2}}^{T_{1}-1}L(T_{1}+i,(\theta+1)T_{1}+T_{2}+i).$
The index function of modality 1 then becomes
$\gamma_{1}(\theta)=\min_{k}\frac{C_{a}(\theta,k)+C_{b}(\theta,k)}{kT_{1}},$
where $C_{a}(\theta,k)$ quantifies the total future cost, while $C_{b}(\theta,k)$ adjusts this cost according to different transmission times.
@@
Due to the non-monotonic AoI functions, our index function is not necessarily monotonic. This result generalizes the two-source scheduling problem in remote estimation [23], where the estimation error is a monotonic function of AoI. Specifically, when the inference error is a non-decreasing function of AoI and $T_{1}=T_{2}=1$, the index function of modality 1 reduces to
$\gamma_{1}(\theta)=\min_{k}\frac{\sum_{j=\theta+2}^{\theta+k+1}L(1,j+1)}{k}=L(1,\theta+3),$
where the last equality holds as the minimum is achieved when $k=1$ since $L$ is non-decreasing. Then, the optimal policy for modality 1 is $\tau_{1}^{*}=\min\{\min\{\theta:L(1,\theta+3)\geq\bar{L}_{\mathrm{opt}}\},\tau_{\max}\}$, which is equivalent to the result in [23, Proposition 3.5].
@@
A similar index-based threshold policy has been studied in the single-source remote inference setting [1, 13]. Specifically, under stationary random transmission time, the optimal time to send a new feature is when an AoI-based index function exceeds a threshold. Different from the single-source setting, we demonstrate how to design an index-based threshold policy when the modalities are non-separable.
@@ V Numerical Case Study: Robot State Prediction
Figure 4: Inference error vs. AoI
@@
Figure 5: Performance comparison under varying transmission times
V-A Experimental Setup
Consider a predictor that aims to continuously track the state (e.g., pose and velocity) of a remote robot. The robot gathers both environmental and self-state information using multimodal sensors, such as LiDAR, cameras, and onboard sensors. The predictor infers the robot’s state based on features transmitted from the robot. Due to the high dimensionality and varying sizes of features, transmissions often span multiple time slots and differ across modalities.
We use the OpenAI Bipedal Walker as our robot task, where a four-joint robot must run over stumps, pitfalls, and other obstacles (see [24] for details). The reinforcement learning algorithm used to control the robot is TD3-FORK [25], which achieves state-of-the-art performance on this task. After training the control model, we generate a time-series dataset in the OpenAI Gymnasium simulation environment, consisting of LiDAR rangefinder measurements, robot state information, and joint control signals.
The inference task is to predict the robot’s velocities using sequential LiDAR measurements and control signals as two distinct modalities. We adopt the Long Short-Term Memory (LSTM) neural network as the predictor model, due to its widespread use and effectiveness in time-series forecasting. The network architecture includes an input layer, a hidden layer with 20 LSTM cells, and a fully connected output layer. We use 80% of the dataset for training. To incorporate AoI into model training, we augment the dataset as follows: given any AoI vector $(\delta_{1},\delta_{2})$, we construct the feature-label pairs $(X_{1,t-\delta_{1}},X_{2,t-\delta_{2}};Y_{t})$ for all time index $t$, where the input features are aligned with their corresponding AoI values. The LSTM network takes the AoI vector as two additional input features and is trained on the augmented training dataset.